OpenStudy (anonymous):

How does 3m^2-12/4m^2+8m X 8m^2+16m/6m-12 turn into m+2?

OpenStudy (mathstudent55):

\( \dfrac{3m^2-12}{4m^2+8m} \times \dfrac{ 8m^2+16m}{6m-12} \) Is this the problem? If so, factor every numerator and denominator.

OpenStudy (anonymous):

Yes, I've tried but I keep getting decimals, I don't understand how the whole thing turns into m+2

OpenStudy (mathstudent55):

Can you factor 3m^2 - 12?

OpenStudy (anonymous):

Yes, into (m+2)(m-2)

OpenStudy (mathstudent55):

Every polynomial you need to factor has only two terms. You need to factor 4 binomials.

OpenStudy (mathstudent55):

To factor a binomial. 1. First try to factor a common term from both terms. 2. Then, check for (a) difference of two squares (b) sum of two cubes (c) difference of two cubes 3. Check every term to make sure it is factored completely.

OpenStudy (mathstudent55):

Let's start with the first numerator. 3m^2 - 12. First factor a common term: 3(m^2 - 4) Now factor the difference of 2 squares: 3(m + 2)(m - 2) So now you have this: \(\dfrac{3(m+2)(m - 2)}{4m^2+8m} \times \dfrac{ 8m^2+16m}{6m-12} \) Now, can you factor the other three binomials?

OpenStudy (mathstudent55):

They are easier to factor than the first one. All you need to do is factor a common factor in each case.

OpenStudy (anonymous):

I'm having trouble with 4m^2+8m

OpenStudy (mathstudent55):

Look at the numbers first. What do 4 and 8 have in common?

OpenStudy (anonymous):

multiples of 4 and 2

OpenStudy (mathstudent55):

Yes, but what is the GCF of 4 and 8?

OpenStudy (anonymous):

8

OpenStudy (mathstudent55):